OpenStudy (anonymous):

Solve this equation. 6j = 30 j = _

Nnesha (nnesha):

6 is multiplying with j to cancel out 6 from left side do opposite of multiply both side

OpenStudy (anonymous):

6j=30 6j/6=30/6 J=5

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

Np aha